# ICODE: Modeling Dynamical Systems with Extrinsic Input Information

### v2

@@ Abstract
Learning models of dynamical systems with external inputs, that may be, for example, nonsmooth or piecewise, is crucial for studying complex phenomena and predicting future state evolution, which is essential for applications such as safety guarantees and decision-making. In this work, we introduce Input Concomitant Neural ODEs (ICODEs), which incorporate precise real-time input information into the learning process of the models, rather than treating the inputs as hidden parameters to be learned. The sufficient conditions to ensure the model’s contraction property are provided to guarantee that system trajectories of the trained model converge to a fixed point, regardless of initial conditions across different training processes. We validate our method through experiments on several representative real dynamics: Single-link robot, DC-to-DC converter, motion dynamics of a rigid body, Rabinovich-Fabrikant equation, Glycolytic-glycogenolytic pathway model, and heat conduction equation. The experimental results demonstrate that our proposed ICODEs efficiently learn the ground truth systems, achieving superior prediction performance under both typical and atypical inputs. This work offers a valuable class of neural ODE models for understanding physical systems with explicit external input information, with potential promising applications in fields such as physics and robotics.
@@
Learning, dynamical systems, inputs, ICODE, physical dynamics, neural ODE.
@@ I Introduction
Neural CDEs and ANODEs vs ICODEs: Neural CDEs [14] (for brevity, we refer to them as ”CDEs” throughout the remainder of this paper) are also designed to incorporate incoming information, while Augmented Neural ODEs (ANODEs) [6] are effective in scenarios where state evolution is less critical. However, many real-world models involve complex external environments (such as atypical, for example, piecewise inputs) and couplings of parameters that are challenging to address accurately with existing NODE and its variants, including CDEs and ANODEs. Our experiments demonstrate that ICODEs are better suited for capturing the natural dynamics of systems with complex structural and environmental characteristics. They can be endowed with a significant contraction property despite their intricate structure.
@@ II Neural ODEs with External Input
Differentiating both sides of the transformation $\delta y=L(x)\delta x$, one can obtain
@@
holds true for some $c>0$ and all $x,u$, where $\lambda_{\max}$ denotes the maximal eigenvalue of a matrix.
@@ III Experimental Results on Learning and Prediction of Physical Systems
In this section, we experimentally analyze the performance of ICODEs, compared to CDE [14], NODEs [5], and ANODEs [6]. We select several important physical systems in various fields, to test the generality of the proposed ICODEs. We trained the models on time series data and used them to predict the evolution of the future state with known input. The models’ training and prediction performance was evaluated. All experiments were conducted with an NVIDIA GeForce RTX 3080ti GPU under Pytorch 1.11.0, Python 3.8 (Ubuntu 20.04), and Cuda 11.3.
@@ III-A Single-link Robot
Figs. 2a, 2b show the loss curves of the selected comparative models, under the third input condition ($\Delta u=0.5$). Further statistical details are in Section IV-A. Our method performs the best and CDE behaves the worst in the setting. This mainly results from the control torque not affecting the system through its derivatives, thus, the derivatives become an interference that affects the performance of CDE.
@@ III-B DC-to-DC Converter
ii) It is noteworthy that in the second scenario, the signal of control input $u$ jumps at $0.1s$, $0.4s$, and $0.8s$ (see Fig. 4f for details). This input is more complex, particularly with a switch occurring at $0.8s$, which falls within the testing phase. Neglecting this information could significantly impair prediction accuracy. As shown in Figs. 4d, 4e, ICODE consistently maintains strong performance in both learning and prediction. In contrast, other methods fail to further reduce prediction errors beyond 400 epochs, with continued training leading to severe overfitting.
@@ III-C Motion of Rigid Body
Figure 5: Comparison of training and prediction errors at each time step on the test set for different models on the task of a rigid body.
@@ III-E Glycolytic-glycogenolytic Pathway Model
III-F Heat Conduction Equation
@@ III-F2 Two-Dimensional Case
IV Experimental Analysis and Settings
IV-A Further Statistical Studies
@@
To further explore the robustness of these methods in noisy environments, we added Gaussian noise of different magnitudes to the training data. The prediction results are drawn in Fig. 12, with numerical details in Table VI, where $p$ indicates the noise level. In Fig. 12, the lines represent the average MSE prediction loss and the shaded areas capture the distribution of loss in these three cases ($p=0.01,0.1,0.2$). This figure illustrates that noises will impact the performance of these methods, increasing the prediction errors and uncertainty (greater variance). However, whichever the case, ICODEs remain the best performer. This demonstrates the superiority of ICODEs in terms of robustness.
@@
TABLE VII: Hyperparameter settings on the six experimental tasks.
Single-link Robot $5\times 10^{-3}$ 100 50 3 10 $[0,1]$
DC-DC Converter $5\times 10^{-4}$ 600 60 3 10 $[0,1]$
Rigid Body $1\times 10^{-3}$ 200 60 2 10 $[0,1]$
R-F Equation $5\times 10^{-4}$ 800 60 3 40 $[0,1]$
Glycolytic-glycogenolytic Pathway $5\times 10^{-4}$ 500 60 3 10 $[0,2]$
Heat Conduction (PDE) $2\times 10^{-3}$ 400 200 3 5 $[0,0.018]$
IV-B Hyperparameter Settings in Tasks
In this section, we provide a detailed description of the hyperparameters employed in the experiments, as summarized in Table VII. All methods are evaluated under identical experimental conditions to ensure a fair comparison. Specifically, ”Hidden Layers” refers to the number of hidden layers in the neural networks, while ”Network Width” indicates the number of neurons in each layer. The ”Trajectories” column specifies the number of trajectories used for model training, and the ”Time Interval” represents the temporal length of the trajectories. For the first four experiments, the ANODE model is augmented with 2 dimensions. In the subsequent experiments, the ANODE is augmented with 200 dimensions to deal with the increased complexity of the PDE case. To generate the initial augmented states, we utilize a multi-layer perceptron (MLP) with a single hidden layer. Model evaluation was conducted every 10 epochs, and the model with the lowest MSE was retained.
@@ V-A Input Span
To examine the significant role of the input signal $u$, we conduct an experimental study comparing the performance of different methods under increasing values of $u$. Using the single-link robot and rigid body models as tasks, we vary the value of the control term $u$ within the range $[-k_{u},k_{u}]$. When $k_{u}$ equals 0, there is no input to the system, while larger values of $k_{u}$ indicate greater external influences on the dynamical system. Table VIII presents the training results for various $k_{u}$, where RMSE stands for Root Mean Square Error. When $k_{u}=0$, the performance of our method reduces to that of the standard NODE, showing no superiority over other methods. With $k_{u}=0.2$, the model experiences a small external influence, leading to larger errors in RMSE and MAE for NODE and ANODE, as they lack mechanisms to handle this variation. In this scenario, our method begins to demonstrate advantages in learning and prediction. When $k_{u}=1$, indicating a conspicuous external influence, the prediction error of our method is markedly lower than that of other methods.
TABLE VIII: Performance comparison of different methods in the three scenarios of input span.
@@ V-B Piecewise Inputs
TABLE IX: Performance comparison between CDE and ICODE under various piecewise inputs.
@@
Both ICODEs and CDEs possess the capability to integrate incoming input information. However, our method demonstrates superior performance, as shown in the previous tasks. In this section, we perform a detailed investigation of their differences by examining three types of piecewise input on the task of the single-link robot. These scenarios involve inputs $u_{1},u_{2},u_{3}$ with varying shapes, as illustrated in Fig. 13a. The values of these inputs range from 00 to $1$. The differences among them are characterized by the length of switching, denoted by $\Delta t$. Specifically, $u_{1}$, the smoothest input, transitions from 00 to $1$ over $0.1s$; $u_{2}$ and $u_{3}$ transition over $0.5s$ and $0.01s$, respectively. It is evident that the input curves become steeper as $\Delta t$ decreases. The learning results are presented in Table IX.
@@ VI Model Scaling Experiments
To further investigate the performance of ICODE in a more complex environment, the Glycolytic-glycogenolytic pathway model is employed as the test scenario. Experiments are conducted using various network architectures, including $2-$ and $3-$layer configurations, with network widths ranging from $4$ to $32$. The results are presented in Figure 14. The training loss, represented by blue lines with the corresponding Y-axis on the left (blue-colored axis), and the validation loss, represented by red lines with the corresponding Y-axis on the right (red-colored axis), are shown. For a fixed number of layers, darker lines correspond to more complex network structures. The figure demonstrates that increasing the network scale effectively reduces training loss and achieves lower validation loss, resulting in more accurate predictions. These findings highlight a scaling law in ICODE, suggesting that problems of varying complexity can be addressed by appropriately adjusting the network structure.

### v4

@@ Abstract
Learning models of dynamical systems with external inputs, which may be, for example, nonsmooth or piecewise, is crucial for studying complex phenomena and predicting future state evolution, which is essential for applications such as safety guarantees and decision-making. In this work, we introduce Input Concomitant Neural ODEs (ICODEs), which incorporate precise real-time input information into the learning process of the models, rather than treating the inputs as hidden parameters to be learned. The sufficient conditions to ensure the model’s contraction property are provided to guarantee that system trajectories of the trained model converge to a fixed point, regardless of initial conditions across different training processes. We validate our method through experiments on several representative real dynamics: Single-link robot, DC-to-DC converter, motion dynamics of a rigid body, Rabinovich-Fabrikant equation, Glycolytic-glycogenolytic pathway model, and heat conduction equation. The experimental results demonstrate that our proposed ICODEs efficiently learn the ground truth systems, achieving superior prediction performance under both typical and atypical inputs. This work offers a valuable class of neural ODE models for understanding physical systems with explicit external input information, with potentially promising applications in fields such as physics and robotics. Our code is available online at https://github.com/EEE-ai59/ICODE.git.
@@
Learning, dynamical systems, ICODE, physical dynamics, neural ODE.
@@ I Introduction
Neural CDEs and ANODEs vs ICODEs: Neural CDEs [14] (for brevity, we refer to them as ”CDEs” throughout the remainder of this paper) are also designed to incorporate incoming information, while Augmented Neural ODEs (ANODEs) [6] are effective in scenarios where state evolution is less critical. A brief introduction to these models and their distinctions from ICODE can be found in appendix A. However, many real-world models involve complex external environments (such as atypical, for example, piecewise inputs) and couplings of parameters that are challenging to address accurately with existing NODE and its variants, including CDEs and ANODEs. Our experiments demonstrate that ICODEs are better suited for capturing the natural dynamics of systems with complex structural and environmental characteristics. They can be endowed with a significant contraction property despite their intricate structure.
@@ II Neural ODEs with External Input
The proof methodology can be found in [27]. First, differentiating both sides of the transformation $\delta y=L(x)\delta x$, one can obtain
@@
holds true for some $c>0$ and all $x,u$, where $\lambda_{\max}$ denotes the maximum eigenvalue of a matrix.
@@ III Experimental Results on Learning and Prediction of Physical Systems
In this section, we experimentally analyze the performance of ICODEs, compared to CDE [14], NODEs [5], and ANODEs [6]. We select several important physical systems in various fields, to test the generality of the proposed ICODEs. We trained the models on time series data and used them to predict the evolution of the future state with known input. The models’ training and prediction performance was evaluated. All experiments were conducted with an NVIDIA GeForce RTX 3080ti GPU under Pytorch 1.11.0, Python 3.8 (Ubuntu 20.04), and Cuda 11.3. The detailed description of the hyperparameter setting can be found in appendix C.
@@ III-A Single-link Robot
Figs. 2a, 2b show the loss curves of the selected comparative models, under the third input condition ($\Delta u=0.5$). Further statistical details are in Section IV. Our method performs the best and CDE behaves the worst in the setting. This mainly results from the control torque not affecting the system through its derivatives, thus, the derivatives become an interference that affects the performance of CDE.
@@ III-B DC-to-DC Converter
ii) It is noteworthy that in the second scenario, the signal of control input $u$ jumps at $0.1s$, $0.5s$, and $0.8s$ (see Fig. 4f for details). This input is more complex, particularly with a switch occurring at $0.8s$, which falls within the testing phase. Neglecting this information could significantly impair prediction accuracy. As shown in Figs. 4d, 4e, ICODE consistently maintains strong performance in both learning and prediction. In contrast, other methods fail to further reduce prediction errors beyond 400 epochs, with continued training leading to severe overfitting.
@@ III-C Motion of Rigid Body
Figure 5: Comparison of prediction errors at each time step on the test set for different models on the task of a rigid body.
@@ III-E Glycolytic-glycogenolytic Pathway Model
III-F Swing Equation
In this example, we analyze the performance of ICODE in multi-agent cooperative systems by considering the swing equation [34], which models the dynamics of synchronous generators or motors in an electrical power system. The dynamics of each node are typically described by the following second-order differential equation:
$M_{i}\frac{d^{2}\theta_{i}}{dt^{2}}+D_{i}\frac{d\theta_{i}}{dt}=P_{i}-\sum_{j=1}^{N}K_{ij}\sin(\theta_{i}-\theta_{j}),$ (11)
where $M_{i}$ represents the moment of inertia of generator $i$, $D_{i}$ is the damping coefficient accounting for the system’s frequency damping effects, and $\theta_{i}$ is the angular position (phase angle) of generator $i$. The term $P_{i}$ denotes the power input to generator $i$, and $K_{ij}$ is the electrical power transmission coefficient between nodes $i$ and $j$, which is related to the electrical conductance and link parameters of the transmission line. The term $\sin(\theta_{i}-\theta_{j})$ represents the power flow between nodes. To convert this system to a first-order differential equation form, we introduce state variables $\omega_{i}:=\frac{d\theta_{i}}{dt}$, representing the angular velocity. Consequently, the state dynamics of each node can be expressed as the following two sets of first-order differential equations: $\frac{d\theta_{i}}{dt}=\omega_{i},\;\frac{d\omega_{i}}{dt}=\frac{1}{M_{i}}\left(P_{i}-\sum_{j=1}^{N}K_{ij}\sin(\theta_{i}-\theta_{j})-D_{i}\omega_{i}\right).$ We consider a network of 10 nodes. The corresponding topology graph is provided in appendix B. The simulation is conducted over the time interval $[0s,5s]$. The external input $P_{i}$ follows the same pattern as depicted in Fig. 4f, varying between 0 and 1 at $t=0.5s,2.5s,$ and $4.5s$. The parameters $M_{i}$ and $D_{i}$ are independently sampled from uniform distributions over the intervals $[0.3,0.9]$ and $[0.7,1.3]$, respectively. A total of 128 trajectories are generated and partitioned into batches, each containing 16 trajectories. In the ANODE framework, we augment the system with 50 additional state variables.
Figure 9: Comparison of training loss and prediction error across time steps for various models applied to the Swing Equation.
The results are presented in Fig. 9. It is clear that ANODE achieves the lowest training loss, which is expected given that the augmented states enhance the network’s degrees of freedom and fitting capacity. However, the prediction error curves indicate that this improved fitting ability does not translate well to generalization. A similar issue is observed with CDE, as its structure does not adapt effectively to these models. In contrast, ICODE consistently demonstrates superior performance in this scenario.
III-G Heat Conduction Equation
@@ III-G2 Two-Dimensional Case
IV Further Statistical Studies
@@
To further explore the robustness of these methods in noisy environments, we added Gaussian noise of different magnitudes to the training data. The prediction results are drawn in Fig. 13, with numerical details in Table VI, where $p$ indicates the noise level. In Fig. 13, the lines represent the average MSE prediction loss, and the shaded areas capture the distribution of loss in these three cases ($p=0.01,0.1,0.2$). This figure illustrates that noises will impact the performance of these methods, increasing the prediction errors and uncertainty (greater variance). However, whichever the case, ICODEs remain the best performer. This demonstrates the superiority of ICODEs in terms of robustness.
We also investigated the scenario in which the measured input $u$ is affected by measurement noise. To simulate this, we introduced Gaussian noise of varying magnitudes into the measured $u$, while ensuring that the system trajectories are generated without these disturbances. Specifically, noise was added at proportions of $10\%$, $20\%$, $40\%$, and $500\%$. The resulting noisy input signals are illustrated in Fig. 14, with the corresponding prediction errors shown in Fig. 15.
As expected, ICODE achieves the best performance when $u$ is noise-free (represented by the blue curve). When a small amount of noise is introduced, the prediction error remains largely unaffected. Under extreme noise conditions ($500\%$), where the external input $u$ carries little useful information, ICODE’s performance degrades to a level comparable to that of NODE. Notably, despite the presence of significant disturbances in the input signal, ICODE remains effective. These results highlight ICODE’s robustness to noise in the input $u$.
@@
Figure 14: The input $u$ with measurement noises of different magnitudes.
Figure 15: Prediction errors on the test set for ICODE under different measurement noises in $u$, compared with NODE.
@@ V-A Input Span
To examine the significant role of the input signal $u$, we conduct an experimental study comparing the performance of different methods under increasing values of $u$. Using the single-link robot and rigid body models as tasks, we vary the value of the control term $u$ within the range $[-k_{u},k_{u}]$. When $k_{u}$ equals 0, there is no input to the system, while larger values of $k_{u}$ indicate greater external influences on the dynamical system. Table VII presents the training results for various $k_{u}$, where RMSE stands for Root Mean Square Error. When $k_{u}=0$, the performance of our method reduces to that of the standard NODE, showing no superiority over other methods. With $k_{u}=0.2$, the model experiences a small external influence, leading to larger errors in RMSE and MAE for NODE and ANODE, as they lack mechanisms to handle this variation. In this scenario, our method begins to demonstrate advantages in learning and prediction. When $k_{u}=1$, indicating a conspicuous external influence, the prediction error of our method is markedly lower than that of other methods.
TABLE VII: Performance comparison of different methods in the three scenarios of input span.
@@ V-B Piecewise Inputs
TABLE VIII: Performance comparison between CDE and ICODE under various piecewise inputs.
@@
Both ICODEs and CDEs possess the capability to integrate incoming input information. However, our method demonstrates superior performance, as shown in the previous tasks. In this section, we perform a detailed investigation of their differences by examining three types of piecewise input on the task of the single-link robot. These scenarios involve inputs $u_{1},u_{2},u_{3}$ with varying shapes, as illustrated in Fig. 16a. The values of these inputs range from 00 to $1$. The differences among them are characterized by the length of switching, denoted by $\Delta t$. Specifically, $u_{1}$, the smoothest input, transitions from 00 to $1$ over $0.1s$; $u_{2}$ and $u_{3}$ transition over $0.5s$ and $0.01s$, respectively. It is evident that the input curves become steeper as $\Delta t$ decreases. The learning results are presented in Table VIII.
@@ VI Model Scaling Experiments
To further investigate the performance of ICODE in a more complex environment, the Glycolytic-glycogenolytic pathway model is employed as the test scenario. Experiments are conducted using various network architectures, including $2-$ and $3-$layer configurations, with network widths ranging from $4$ to $32$. The results are presented in Fig. 17. Blue lines on the left represent the training loss, and red lines on the right represent the validation loss. For a fixed number of layers, darker lines correspond to more complex network structures. In the two subplots, generally, the deeper the color of the line, the lower the loss. This demonstrates that increasing the network scale effectively reduces training loss and achieves lower validation loss, resulting in more accurate predictions. These findings highlight a scaling law in ICODE, suggesting that problems of varying complexity can be addressed by appropriately adjusting the network structure.
@@ VII Conclusion
Appendix
A. A brief introduction of existing neural ODEs and their comparison with ICODE
In this section, we briefly introduce the NODE, CDE, and ANODE and clarify the core distinctions between CDE and ICODE.
1) Neural Ordinary Differential Equations
The Neural Ordinary Differential Equation (NODE or neural ODE) is defined as follows [5]:
$y(0)=y_{0},\quad\frac{\mathrm{d}y}{\mathrm{d}t}(t)=f_{\theta}(t,y(t)),$ (12)
where $y_{0}\in\mathbb{R}^{d_{y}}$ denotes the initial condition, $\theta$ is a vector of learnable parameters, and $f_{\theta}:\mathbb{R}\times\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}^{d_{y}}$ represents a neural network. In many cases, $f_{\theta}$ is implemented using a standard neural network architecture, such as a feedforward or convolutional network.
2) Neural Controlled Differential Equation
The Neural Controlled Differential Equation (CDE) enhances classical NODEs from a mathematical perspective, enabling the effective processing of incoming data.
The state of a neural ODE can be expressed in the form of the following Riemann integral:
$y(0)=y_{0},\quad y(t)=y(0)+\int_{0}^{t}f_{\theta}(s,y(s))\,\mathrm{d}s,$ (13)
which naturally suggests an extension to a Riemann-Stieltjes integral:
$y(0)=y_{0},\quad y(t)=y(0)+\int_{0}^{t}f(y(s))\,\mathrm{d}x(s),$ (14)
where $x$ is a high-dimensional function with respect to time $t$.
A key limitation of ODEs, as given in (12) and (13), is that their solutions are solely determined by the initial condition $y(0)$ and the function $f_{\theta}$, with no direct mechanism for incorporating data that arrives at later time steps. In contrast, in (14), the integrator $x$ is explicitly time-dependent, allowing temporal information to be naturally incorporated into the system. This fundamental distinction underpins the development of CDEs.
We now introduce the formal definition of CDEs:
Let $T>0$ and let $d_{x},d_{y}\in\mathbb{N}$. Let $x\colon[0,T]\to\mathbb{R}^{d_{x}}$ be a continuous function of bounded variation. Let $f\colon\mathbb{R}^{d_{y}}\to\mathbb{R}^{d_{y}\times d_{x}}$ be Lipschitz continuous. Let $y_{0}\in\mathbb{R}^{d_{y}}$. A continuous path $y\colon[0,T]\to\mathbb{R}^{d_{y}}$ is said to solve a controlled differential equation, controlled or driven by $x$, if it satisfies Equation (14) for $t\in(0,T]$. Here, $\mathrm{d}x(s)$ denotes a Riemann-Stieltjes integral, and $f(y(s))\,\mathrm{d}x(s)$ refers to a matrix-vector multiplication.
As can be seen, CDE incorporates incoming data by integrating the known information $x(s)$. Given the differentiable assumption, a CDE is solved by reducing it to an ODE [14, 35]. To be more specific, let
$g_{\theta,x}(y,s)=f_{\theta}(y)\frac{\mathrm{d}x}{\mathrm{d}s}(s),$
so that for $t\in(0,T]$,
$y(t)$ $=y(0)+\int_{0}^{t}f_{\theta}(y(s))\,\mathrm{d}x(s)$
$=y(0)+\int_{0}^{t}f_{\theta}(y(s))\frac{\mathrm{d}x}{\mathrm{d}s}(s)\,\mathrm{d}s$
$=y(0)+\int_{0}^{t}g_{\theta,x}(y(s),s)\,\mathrm{d}s.$
Suppose that we have known information $x$ with $x(t_{i}):=(t_{i},(u_{i})^{\top})^{\top}$. Then we have $g_{\theta,x}(y,s)=f_{\theta}(y)\frac{\mathrm{d}x}{\mathrm{d}s}(s)=[f_{\theta,1}(y)\;f_{\theta,2:}(y)][1,(\frac{\text{d}u}{\text{d}t})^{\top}]^{\top}(s)=f_{\theta,1}(y)+f_{\theta,2:}(y)\frac{\text{d}u}{\text{d}t}(s)$, where $f_{\theta,1}$ and $f_{\theta,2:}$ are neural networks to be trained. This gives
[…]
Compare it with our ICODE:
[…]
where $f$ and $g$ are neural networks to be trained. It can be seen that the difference lies in whether the derivative operation is necessary before incorporating the external information $u$ into the model.
3) Augmented Neural ODE
The Augmented Neural Ordinary Differential Equation (ANODE) has been shown to learn representations that preserve the topology of the input space, resulting in certain functions that the NODE cannot represent. To address this limitation, ANODE [6] augments the space in which the Ordinary Differential Equation (ODE) is solved, enabling the model to leverage the additional dimensions for learning more complex functions. Specifically, the NODE (12) is enhanced to the following form:
$h(0)=\begin{bmatrix}y(0)\\ \mathbf{0}\end{bmatrix}\in\mathbb{R}^{d_{y}+d_{a}},\quad\frac{\mathrm{d}}{\mathrm{d}t}h(t)=\tilde{f}_{\theta}\left(h(t),t\right).$
Here, $d_{a}$ represents the augmented dimensions.
ANODE is especially advantageous when the continuous system includes latent variables or unmodeled dynamics, which are common in dynamical systems with non-unique solutions. However, in the absence of such complexities, ANODE may result in diminished generalization performance due to its excessive flexibility. When the order and state variables of a continuous system are well-defined, incorporating topology preservation as prior information may be disrupted by ANODE. Unlike classification tasks that utilize dynamic systems, our focus is on modeling dynamic systems directly. Consequently, augmenting additional variables is often unnecessary in this context. As demonstrated in the experimental section, while ANODE may exhibit minimal error on the training set, its performance tends to be suboptimal on the test set. significantly weaker.
TABLE IX: Hyperparameter settings on the seven experimental tasks.
Single-link Robot $5\times 10^{-3}$ 100 50 3 10 $[0,1]$
DC-DC Converter $5\times 10^{-4}$ 600 60 3 10 $[0,1]$
Rigid Body $1\times 10^{-3}$ 200 60 2 10 $[0,1]$
R-F Equation $5\times 10^{-4}$ 800 60 3 40 $[0,1]$
Glycolytic-glycogenolytic Pathway $5\times 10^{-4}$ 500 60 3 10 $[0,2]$
Swing Equation (multi-agent) $5\times 10^{-4}$ 200 60 3 128 (batch size 16) $[0,5]$
Heat Conduction (PDE) $2\times 10^{-3}$ 400 200 3 5 $[0,0.018]$
B. Specific details about the experiments
The topology structure of the $10$ nodes in the experiment of the Swing Equation is shown in Fig. 18. The trajectories of these agents (synchronous generators) are presented in Fig. 19.
C. Hyperparameter settings
This section provides a comprehensive description of the hyperparameters used in the experiments, as summarized in Table IX. To ensure a fair comparison, all methods are evaluated under identical experimental conditions. Specifically, Hidden Layers denotes the number of hidden layers in the neural networks, while Network Width represents the number of neurons per layer. The Trajectories column specifies the number of trajectories used for model training, and Time Interval indicates the temporal length of these trajectories.
For the first four experiments, the ANODE model is augmented with $2$ additional dimensions. In the fifth experiment, this augmentation is increased to $10$ dimensions. In subsequent experiments, the ANODE model is expanded to $200$ dimensions to accommodate the increased complexity of the partial differential equation (PDE) case. The initial augmented states are generated using a multi-layer perceptron (MLP) with a single hidden layer.
The implementation employs the Softplus activation function and the Adam optimizer. Model evaluation is performed every $10$ epochs, with the model achieving the lowest mean squared error (MSE) retained for final analysis.
Figure 18: The topology structure of the power system network.
Figure 19: The trajectories of the agents (synchronous generators) in the network.